A spark is generated in an automobile spark plug when there is an electric potential of 3000 V across the electrode gap. If 60 W of power is generated in a single spark that delivers a total charge of 3 nC, how long does it take for the spark to travel across the gap?

50 ns

75 ns

125 ns

150 ns

225 ns
Answer #1

Here,V=3000 volts and power=60 W.

So,60=3000*I

=>I=60/3000=0.02 A.

Now,current=q/t where q is charge and t is time.

Here, q=3 nC=3*10^-9 C.

So, 0.02=(3*10^-9)/time

=>time=3*10^-9/0.02=150*10^-9 s=150 ns.(fourth option)